OpenStudy (anonymous):

solve the quadratic equation by completing the square -3x^2+9x=1

hartnn (hartnn):

what have you tried ? I would have started by dividing the entire equation by -3

OpenStudy (anonymous):

I don't understand how to solve it...

OpenStudy (anonymous):

that's what I need help understanding

hartnn (hartnn):

you need to complete the square on left side like for \(x^2+2bx\) you will add b^2 to make it \(x^2+ 2bx+b^2 \) which indeed is a perfect square : \((x+b)^2\)

hartnn (hartnn):

so for -3x^2 +9x =1 first divide entire equation by -3 to make the co-efficient of x^2 as 1 then we will start completing the square

OpenStudy (anonymous):

so by dividing the whole thing by -3 x^2 becomes 1?

OpenStudy (anonymous):

1x^2 +-3x

hartnn (hartnn):

the 'co-efficient' of x^2 will become 1, right ? as -3x^2/-3 = x^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

soooo, why is it = x^2

hartnn (hartnn):